This year (10 years after you first took out the loan), you check your loan balance. Only part of your payments have been going to pay down the loan; the rest has been going towards interest. You see that you still have $100,400 left to pay on your loan. Your house is now valued at $180,000.


Your current situation


How much of the original loan have you paid off? (i.e, how much have you reduced the loan balance by? Keep in mind that interest is charged each month - it's not part of the loan balance.)

How much money have you paid to the loan company so far (over the last 10 years)?

How much interest have you paid so far (over the last 10 years)?

How much equity do you have in your home (equity is value minus remaining debt)


1
Expert's answer
2021-04-14T14:36:47-0400

(1) mortgage amount=130000

10%down payment=(13000)

Current mortgage=117000

Loan balance=(100400)

Original loan paid =16600

(2)total loan paid=loan paid+interest

"=16600+7583=24183"

(3) interst="Principal\times interest rate\times time"

"=130000\times\frac{0.07}{12}\times 10"

"=7583"

(4) Equity=house value-remaining debt

"180000-100400=79600"